OpenStudy (anonymous):

5. Solve the following Gas Law Problems. Be sure to include which law applies in the solution. a) A gas has a volume of 100 ml when the pressure is 735 mm. What volume will the gas occupy at 700 mm? b) Change the following temperatures to Kelvin scale: 1) 20 C 2) 85 C 3) - 15 C 4) - 190 C c) Given 90 ml of Hydrogen gas collected with the temperature is 27 degrees C. What volume will the Hydrogen occupy at 42 C? d) A gas measures 500 ml at a temperature of -23 C. Find its volume at 23 C. e)

OpenStudy (rogue):

5a. You can use boyle's law to find the new volume (V2), everything else is given.Also, by mm, the question probably means mmHg (milimeters of Mercury; this is the same thing as torr. \[P_1 V_1 = P_2 V_2 \rightarrow \frac {P_1}{V_2} = \frac {P2}{V_1}\]5b. Just use the conversion formula for Celsius to Kelvin. \[k = C + 273.15\]So for example, 0 C would be 273.15 k or 100 C would be 373.15 k. 5c. Use Charles's law to find T2.\[V_1 T_2 = V_2 T_1 \rightarrow \frac {V_1}{T_1} = \frac {V_2}{T_2}\]5d. Use Charles's law to find V2.